OpenStudy (anonymous):

What is the vertex for the graph of y = 2x2 + 12x + 10

OpenStudy (anonymous):

y'=4x+6=0 x=-1.5 y=-3.5

OpenStudy (anonymous):

its looking for vertex...

OpenStudy (anonymous):

(-1.5,-3.5) yes

OpenStudy (anonymous):

well shouldnt it be multiplied by 2? and thats none of the answer choices

OpenStudy (anonymous):

sorry 4x+12=0 x=-3 y=4

OpenStudy (anonymous):

I thought that was 6 instead 12

OpenStudy (anonymous):

the closes answer on here is (-3,8)....

OpenStudy (anonymous):

I guess you changed it!!didn't it?

OpenStudy (anonymous):

2*9-12*3+10=4

OpenStudy (anonymous):

i didnt change anything?! haha

OpenStudy (anonymous):

Ok !!I don't no where is my mind!!

OpenStudy (anonymous):

(-3,4)is vertex

OpenStudy (anonymous):

(−3, −8) (−3, −44) (3, 64) (3, 58)

OpenStudy (anonymous):

sir thats not an answer! lol

OpenStudy (anonymous):

put -3 in equation so get 4 !!

OpenStudy (anonymous):

yes I mistake 18-36+10=-8

OpenStudy (anonymous):

so its (-3,-8)

OpenStudy (anonymous):

yes